# LLM-Agent-UMF: LLM-based Agent Unified Modeling Framework for Seamless Integration of Multi Active/Passive Core-Agents

### v2

@@ 3.3 Active/Passive Core-Agent Classification
3.3.1 Active Core-Agents
Active core-agents encompass all five modules described in Section 3.2 and illustrated in Figure 3, but what differentiates an active from a passive core-agent is its managerial aspect. An active core-agent is characterized by its leading position in the agent as the orchestrator of other components, so naturally, it requires a planning module to divide tasks into subtasks and collaborates with the memory module to provide the necessary context, analyze information, and make decisions. Consequently, we consider an active core-agent to be stateful, meaning it can maintains information about its past interactions and states over time. This is facilitated by an adaptive memory that captures and stores various aspects of the agent’s lifecycle, allowing it to use this historical data to inform future actions and decisions. The profile module role is emphasized in the active core-agent category, because it guides the LLM’s behavior in a certain direction. Furthermore, the security module plays a prominent role in safeguarding the communication between the LLM and the human, ensuring a reliable exchange; Acting as an intermediary, the core-agent safeguards the LLM from threats such as jailbreak attempts and protects user data privacy by implementing safety measures as outlined in Table 1.
@@ 3.3.1 Active Core-Agents
Modules Sub-modules / Methods Active Core-Agent Passive Core-Agent
@@
Throughout our research on the state of the art, we observed that LLM-based agents are recently built upon active core-agents performing tasks from planning to execution [5]. As highlighted in [18], active core-agents are more effective because they incorporate planning and memory modules, which enable them to reason, plan, and execute tasks efficiently. This structure allows the agent to adapt to changing situations and make informed decisions, making the system more robust and capable.
However, relying solely on active core-agents would increase the complexity of the agent, which can lead to scalability issues and negatively impacts the maintainability of the agent as it will hinder and complicate future improvement efforts. As noted in [18], "the complexity of the agent system grows exponentially with the number of tasks it needs to perform". Therefore, rather than centralizing responsibilities on one entity, it would be more beneficial and adhering to the Single Responsibility Principle, if we leverage other core-agents to granulate task execution and reduce the complexity of the agent system. This approach is supported by the concept of "separation of concerns" in software engineering, which emphasizes the importance of dividing responsibilities among multiple components to improve system modularity and maintainability. By distributing tasks among multiple core-agents, we can reduce the cognitive load on individual core-agents, improve system efficiency, and enhance overall performance.
3.3.2 Passive Core-Agents
@@ 3.4.1 Uniform multi-core agent
  • Uniform active core-agents architecture deals with the interaction of a group of active core-agents in one system as illustrated in Figure 11. As opposite to passive core-agents that operate solely with action and security modules, active core-agents possess all five modules (Planning, Memory, Profile, Action and Security) enabling them to manage complex cognitive tasks. This architecture may be seen as a better alternative to uniform multi-passive core-agents’ design due to its wider range of capabilities and functionalities. However, due to the authoritative nature of the active entities, the multi-active core-agent design is more complex than the one exclusively based on passive core-agents. The inclusion of multiple active elements in one system introduces challenges similar to those in multi-agent systems. For instance, effective communication among active core-agents is paramount; given their dynamic nature, timely and accurate exchange of information _such as inter-sibling core-agent feedback and status updates_ is crucial to ensure cohesive operation of the agent. As the number of active core-agents grows, managing intra-communication becomes increasingly complex resulting in frequently emerging synchronization issues. This is why multi-active core-agent systems potentially necessitate consensus algorithms, such as Raft [61, 62], to elect a leader.
@@ 3.4.2 Hybrid multi-core agent
To leverage the strength of both passive and active core-agents architectures, we propose an optimal system depicted in Figure 12. It integrates one active entity as the manager with multiple passive entities functioning as workers within a unified system. One managerial aspect of the active core-agent is its ability to configure dynamically the profile of LLMs, enabling them to effectively utilize passive core-agents for handling specific tasks. This configuration leverages the parallel execution capabilities of numerous passive core-agents under the guidance and leadership of an active core-agent, empowering the system to handle wider range of tasks, while preserving a comfortable level of flexibility, extendibility and scalability.
@@
In conclusion, the modularity of core-agents contributes to guaranteeing the composability within the agent architecture from a software perspective. It facilitates the seamless integration of new passive core-agents within a single agent system as the system scales, obviating the need for a transition to a multi-agent system. Furthermore, this architecture tackles scalability and adaptability challenges by adhering to the Open/Closed Principle (OCP), enhancing core-agents’ integration across evolving systems, and fostering robustness and flexibility. As outlined in the paper [7], the expansion of the system necessitates dynamic scaling to accommodate growing demands and ensure optimal performance. This entails adaptive capabilities such as increasing the number of agents or utilizing larger LLMs. These challenges are effectively addressed by architectures based on multiple core-agents primarily due to the unitary role a core-agent plays within the agent system. In fact, our framework allows the active core-agent to dynamically incorporate or detach passive core-agents, as illustrated by the switch linking the leader active core-agent and the passive core-agent (2) in Figure 13. In the following section, we discuss the results of our work.
4 Results and discussion
LLM-based agents have traditionally been discussed and conceptualized as intricate systems with different inner functional entities, such as LLMs and complementary software components, treated in an intertwined manner. Our proposed LLM-based Agent Unified Modeling Framework (LLM-Agent-UMF) was shaped to overcome these issues and promote a clear delineation of components and responsibilities. In this section, we present and analyze the outcomes derived from adopting the LLM-Agent-UMF through diverse evaluation approaches. Firstly, Section 4.1 evaluates the newly introduced core-agent terminology by applying the LLM-Agent-UMF to existing systems that do not explicitly identify themselves as agents. Secondly, Section 4.2 dives deeper into understanding and dissecting the internal structure of thirteen LLM-based agents by leveraging our proposed framework. This is performed not only on open-source agents but also on proprietary ones by formulating hypothesis based on their observed behavior. Through this process, we classify core-agents into active and passive types and identify critical modules that may have been overlooked or underemphasized during development. Finally, Section 4.3 presents five novel architectures based on the LLM-Agent-UMF to model multi-core agents by combining characteristics from separate agents not intended for fusion in their original design.
@@ 4.1 Evaluation of the new core-agent terminology
One of the main reasons why agent definition remains misunderstood lies in the disparity between how researchers and practitioners use terminology to describe their work. For instance, some studies might not explicitly refer to their LLM-based system as an "agent", even though it exhibits characteristics commonly associated with agents such as autonomy, adaptability, and goal-directed behavior. The ToolLLM [59] research paper provides a prime example of this discrepancy. While the authors indeed utilize API retriever component in conjunction with their LLM, they refrain from using the term "agent". The same can be observed in the Toolformer [38] work, which implicitly incorporates software components to execute API call requests generated by LLMs but similarly avoids referencing the system as an agent.
@@
Based on the definition that we established for the core-agent component, we successfully identified core-agents within multiple LLM-based systems, mainly Toolformer and ToolLLM. On the one hand, Toolformer incorporates a simple-structured core-agent that includes only the action module responsible for API execution. On the other hand, ToolLLM adopted a more sophisticated approach. In fact, we recognize the use of a core-agent encompassing two modules: an action module represented by the neural API retriever and an implicate planning module responsible of managing the flow of information within the agent: The core-agent intercept user instructions, leverage the API retriever to gather relevant APIs, relays the APIs to the LLM for response generation, executes the requested APIs, and finally returns the outcome back to the LLM for final user response formulation. Thus, the identification of core-agents in these systems, accompanied with the presence of LLMs elucidates that Toolformer and ToolLLM are indeed LLM-based agents.
@@ 4.2 Evaluation of active/passive core-agent internal structure delineation
Going further in our analysis, we study ToolLLM [59] and Gorilla [70], and we examine the latter two modes: "zero-shot" and "with retriever". In the zero-shot mode, Gorilla behaves similarly to the three previously discussed agents by acting as an LLM-based agent with a passive core-agent that possesses an action module for API call execution. However, in the second mode, functioning similarly to ToolLLM, Gorilla utilizes an API retriever to gather API recommendations and fetch documentation. This process necessitates a software component for orchestrating interactions with LLMs and tools; this component, which we label as the core-agent, was not highlighted in the original papers. It is inherent that this core-agent includes a planning module incorporated as the API retriever and requires a minimal memory module to manage the internal state of the agent in case of multi-step instructions.
We complement our analysis by deeply evaluating the case of GPT4Tools [36], which took inspiration from its predecessors and focused on finetuning the LLM to utilize tools but extended its capabilities to leverage visual models such as image segmentation and generation models. Once trained using their technique, the LLM is capable of responding to instruction by performing multiple steps to achieve a targeted goal. To satisfy the needs of such an LLM, a software component must be integrated to listen to its requests, execute actions and either return textural responses or store intermediate outputs such as images in its memory to be utilized in subsequent steps. This software entity, identified as the core-agent, lacks any established authoritative nature but requires basic planning and memory modules to manage multi-step subtasks. Therefore, we classify it as an active core-agent.
@@
  • Hypothesis 2: Security measures are implemented independently outside the scope of the main LLM. The input is handled before being forwarded to the LLM, and the output is monitored after its generation and before being presented to the user. According to this flow of events, we observe a minimum level of algorithmic planning thus the need for an active core-agent responsible for managing the guardrails of ChatGPT. However, such simple planning designed for a specific goal _ensuring safeguarding workflow_ does not require a memory module nor a profile module.
@@
Through this exercise, we demonstrate how the LLM-Agent-UMF can assist developers in reevaluating their agent designs and potentially improving upon them. Specifically, out of all studied agents that utilize external tools, 78% (7 over 9) did not incorporate necessary security measures to handle privacy concerns. Only ChemRow and ChatGPT demonstrated an ability to manage this aspect effectively. ChemRow utilized IBM Research’s tools, which fundamentally emphasize privacy protection [72], while ChatGPT takes an additional step by requiring user validation before sharing any data with third-parties. This highlights the critical importance of addressing privacy concerns and safeguarding against information leakage to ensure robustness and trustworthiness in AI-powered systems.
@@ 4.3.2 ToolLLM and ChatDB as a multi active core-agent system
As mentioned in Section 3.4.2, incorporating multiple active core-agents within one system could pose challenges. To evaluate this scenario, two architectural variants were explored: LA2-A and LA2-B. In LA2-A, Figure 16, both ToolLLM and ChatDB retained their individual functionalities as distinct active core-agents. This case requires synchronization between the two active core-agents and opting for a consensus algorithm like Raft [61] would be a knowledgeable choice. To further optimize the memory footprint of the system, there will be one unique instance of Llama 3.1 8B shared between the two active core-agents and each one of them must inject the adequate profile dynamically either as a pluggable trained module like LoRA or using a system prompt.
@@ 4.3.4 Hybrid multi active/passive core-agents system
Nonetheless, it is crucial to address the security vulnerability uncovered through the application of LLM-Agent-UMF within LA4’s design. Indeed, the external API calling mechanisms utilized by Toolformer and Confucius passive core-agents are not monitored or safeguarded against potential information leakage. Following the guidance provided by LLM-Agent-UMF, this issue can be resolved either by implementing a dedicated security module on each of these passive core-agents or as supplementary measures within the active core-agent’s security module, thus centralizing the management of information safeguarding processes. The optimal selection of technologies and implementation details will be explored in future work.
@@ 5 Conclusion and future work
In this paper, we introduce a structural component within LLM-based agents named the "core-agent". This component is engineered to address the architectural ambiguities that software developers encounter and foster better understanding of the interacting entities within LLM-powered agents. We propose the LLM-Agent-UMF, a comprehensive framework for modeling the structure of the agent, explicating each of core-agent’s five modules: planning, memory, profile, action, and security. Subsequently, we classified core-agents into passive and active categories and highlighted their structural and functional differences. Based on this classification, we designed uniform and hybrid multi-core agent architectures. Most prominently, the one-active-many-passive architecture exploits the full potential of both active and passive core-agents, striking a balance between easiness of development and the power of hybrid architectures. By applying our framework to state-of-the-art agents, we identified within their structures core-agents and their constituting internal modules which assisted us in the classification process. This allowed us to recognize the individual characteristics of each agent, discern their limitations and discover potential prospects of merging different functionalities into a single multi-core agent.
Our work prepares the foundation for the development of LLM-based agents with a clear delineated structure that leverages the power of core-agents. The progressive adoption of LLM-Agent-UMF will further attest to its efficiency. An interesting future direction to improve it involves finding solutions to simplify the implementation of multi-active core-agent architectures. In fact, they suffer from challenges related to synchronization that necessitate further investigation. In this context, we see two promising avenues: Integrating a consensus algorithms like Raft to elect a leader responsible for the coordination; Otherwise, integrating a central gateway that is solely responsible of selecting the most suitable active core-agent to handle the user request based on factors like load, availability, and domain. Each active core-agent shall register with the gateway, providing information about their capabilities and status. The selected core-agent processes the task and sends the response back to the user through the gateway.

### v3

@@ 4.3 Active/Passive Core-Agent Classification
4.3.1 Active Core-Agents
Active core-agents encompass all five modules described in Section 4.2 and illustrated in Figure 3, but what differentiates an active from a passive core-agent is its managerial aspect. An active core-agent is characterized by its leading position in the agent as the orchestrator of other components, so naturally, it requires a planning module to divide tasks into subtasks and collaborates with the memory module to provide the necessary context, analyze information, and make decisions. Consequently, we consider an active core-agent to be stateful, meaning it can maintain information about its past interactions and states over time. This is facilitated by an adaptive memory that captures and stores various aspects of the agent’s lifecycle, allowing it to use this historical data to inform future actions and decisions. The profile module role is emphasized in the active core-agent category, because it guides the LLM’s behavior in a certain direction. Furthermore, the security module plays a prominent role in safeguarding the communication between the LLM and the human, ensuring a reliable exchange; Acting as an intermediary, the core-agent safeguards the LLM from threats such as jailbreak attempts and protects user data privacy by implementing safety measures as outlined in Table 1.
As highlighted in [7], active core-agents are more effective because they incorporate planning and memory modules, which enable them to reason, plan, and execute tasks efficiently. This structure allows the agent to adapt to changing situations and make informed decisions, making the system more robust and capable.
However, relying solely on active core-agents would increase the complexity of the agent, which can lead to Scalability issues and negatively impacts the Maintainability of the agent as it will hinder and complicate future improvement efforts. As noted by [7], “the complexity of the agent system grows exponentially with the number of tasks it needs to perform”. Therefore, rather than centralizing responsibilities on one entity, it would be more beneficial and adhering to the Single Responsibility Principle, if we leverage other core-agents to granulate task execution and reduce the complexity of the agent system. This approach is supported by the concept of Separation of Concerns in software engineering, which emphasizes the importance of dividing responsibilities among multiple components to improve system Modularity and Maintainability. By distributing tasks among multiple core-agents, we can reduce the cognitive load on individual core-agents, improve system efficiency, and enhance overall performance.
4.3.2 Passive Core-Agents
@@
Modules Sub-modules / Methods
Active
Core-Agent
Passive
Core-Agent
@@ 4.4.1 Uniform multi-core agent
  • Uniform active core-agents architecture deals with the interaction of a group of active core-agents in one system as illustrated in Figure 11. As opposite to passive core-agents that operate solely with action and security modules, active core-agents possess all five modules (Planning, Memory, Profile, Action and Security) enabling them to manage complex cognitive tasks. This architecture may be seen as a better alternative to uniform multi-passive core-agents’ design due to its wider range of capabilities and functionalities. However, due to the authoritative nature of the active entities, the multi-active core-agent design is more complex than the one exclusively based on passive core-agents. The inclusion of multiple active elements in one system introduces challenges similar to those in multi-agent systems. For instance, effective communication among active core-agents is paramount; given their dynamic nature, timely and accurate exchange of information (such as inter-sibling core-agent feedback and status updates) is crucial to ensure cohesive operation of the agent. As the number of active core-agents grows, managing intra-communication becomes increasingly complex resulting in frequently emerging synchronization issues. This is why multi-active core-agent systems potentially necessitate consensus algorithms, such as Raft [70, 71], to elect a leader.
@@ 4.4.2 Hybrid multi-core agent
To leverage the strength of both passive and active core-agents architectures, we propose an optimal system depicted in Figure 12. It integrates one active entity as the manager with multiple passive entities functioning as workers within a unified system. One managerial aspect of the active core-agent is its ability to configure dynamically the profile of LLMs, enabling them to effectively utilize passive core-agents for handling specific tasks. This configuration leverages the parallel execution capabilities of numerous passive core-agents under the guidance and leadership of an active core-agent, empowering the system to handle wider range of tasks, while preserving a comfortable level of flexibility, extensibility and scalability.
@@
In conclusion, the modularity of core-agents contributes to guaranteeing the composability within the agent architecture from a software perspective. It facilitates the seamless integration of new passive core-agents within a single agent system as the system scales, obviating the need for a transition to a multi-agent system. Furthermore, this architecture tackles Scalability and Flexibility challenges by adhering to the Open-Closed Principle (OCP), enhancing core-agents’ integration across evolving systems, and fostering robustness and flexibility. As outlined in the paper [7], the expansion of the system necessitates dynamic scaling to accommodate growing demands and ensure optimal performance. This entails adaptive capabilities such as increasing the number of agents or utilizing larger LLMs. These challenges are effectively addressed by architectures based on multiple core-agents primarily due to the unitary role a core-agent plays within the agent system. In fact, our framework allows the active core-agent to dynamically incorporate or detach passive core-agents, as illustrated by the switch linking the leader active core-agent and the passive core-agent (2) in Figure 13. In the following section, we evaluate and discuss the results of our work.
5 Evaluation and Discussion
LLM-based agents have traditionally been discussed and conceptualized as intricate systems with different inner functional entities, such as LLMs and complementary software components, treated in an intertwined manner. Our proposed LLM-based Agent Unified Modeling Framework (LLM-Agent-UMF) was shaped to overcome these issues and promote a clear delineation of components and responsibilities. In this section, we present and analyze the outcomes derived from adopting the LLM-Agent-UMF through diverse evaluation approaches. Firstly, Section 5.1 evaluates the newly introduced core-agent terminology by applying the LLM-Agent-UMF to existing systems that do not explicitly identify themselves as agents. Secondly, Section 5.2 dives deeper into understanding and dissecting the internal structure of thirteen LLM-based agents by leveraging our proposed framework. This is performed not only on open-source agents but also on proprietary ones by formulating hypotheses based on their observed behavior. Through this process, we classify core-agents into active and passive types and identify critical modules that may have been overlooked or underemphasized during development. Finally, Section 5.3 presents five novel architectures based on the LLM-Agent-UMF to model multi-core agents by combining characteristics from separate agents not intended for fusion in their original design.
@@ 5.1 Evaluation of the new core-agent terminology
One of the main reasons why agent definition remains misunderstood lies in the disparity between how researchers and practitioners use terminology to describe their work. For instance, some studies might not explicitly refer to their LLM-based system as an “agent”, even though it exhibits characteristics commonly associated with agents such as autonomy, adaptability, and goal-directed behavior. The ToolLLM [42] research paper provides a prime example of this discrepancy. While the authors indeed utilize API retriever component in conjunction with their LLM, they refrain from using the term “agent”. The same can be observed in the Toolformer [38] work, which implicitly incorporates software components to execute API call requests generated by LLMs but similarly avoids referencing the system as an agent.
@@
Based on the definition that we established for the core-agent component, we successfully identified core-agents within multiple LLM-based systems, mainly Toolformer and ToolLLM. On the one hand, Toolformer incorporates a simple-structured core-agent that includes only the action module responsible for API execution. On the other hand, ToolLLM adopted a more sophisticated approach. In fact, we recognize the use of a core-agent encompassing two modules: an action module represented by the neural API retriever and an implicit planning module responsible of managing the flow of information within the agent: The core-agent intercept user instructions, leverage the API retriever to gather relevant APIs, relays the APIs to the LLM for response generation, executes the requested APIs, and finally returns the outcome back to the LLM for final user response formulation. Thus, the identification of core-agents in these systems, accompanied with the presence of LLMs elucidates that Toolformer and ToolLLM are indeed LLM-based agents.
@@ 5.2 Evaluation of active/passive core-agent internal structure delineation
Going further in our analysis, we study ToolLLM [42] and Gorilla [41], and we examine the latter two modes: “zero-shot” and “with retriever”. In the zero-shot mode, Gorilla behaves similarly to the three previously discussed agents by acting as an LLM-based agent with a passive core-agent that possesses an action module for API call execution. However, in the second mode, functioning similarly to ToolLLM, Gorilla utilizes an API retriever to gather API recommendations and fetch documentation. This process necessitates a software component for orchestrating interactions with LLMs and tools; this component, which we label as the core-agent, was not highlighted in the original papers. It is inherent that this core-agent includes a planning module incorporated as the API retriever and requires a minimal memory module to manage the internal state of the agent in case of multi-step instructions.
We complement our analysis by deeply evaluating the case of GPT4Tools [43], which took inspiration from its predecessors and focused on finetuning the LLM to utilize tools but extended its capabilities to leverage visual models such as image segmentation and generation models. Once trained using their technique, the LLM is capable of responding to instruction by performing multiple steps to achieve a targeted goal. To satisfy the needs of such an LLM, a software component must be integrated to listen to its requests, execute actions and either return textual responses or store intermediate outputs such as images in its memory to be utilized in subsequent steps. This software entity, identified as the core-agent, lacks any established authoritative nature but requires basic planning and memory modules to manage multi-step subtasks. Therefore, we classify it as an active core-agent.
@@
  • Hypothesis 2: Security measures are implemented independently outside the scope of the main LLM. The input is handled before being forwarded to the LLM, and the output is monitored after its generation and before being presented to the user. According to this flow of events, we observe a minimum level of algorithmic planning thus the need for an active core-agent responsible for managing the guardrails of ChatGPT. However, such simple planning designed for a specific goal, ensuring safeguarding workflow, does not require a memory module nor a profile module.
@@
Through this exercise, we demonstrate how the LLM-Agent-UMF can assist developers in reevaluating their agent designs and potentially improving upon them. Specifically, out of all studied agents that utilize external tools, 78% (7 out of 9) did not incorporate necessary security measures to handle privacy concerns. Only ChemCrow and ChatGPT demonstrated an ability to manage this aspect effectively. ChemCrow utilized IBM Research’s tools, which fundamentally emphasize privacy protection [76], while ChatGPT takes an additional step by requiring user validation before sharing any data with third-parties. This highlights the critical importance of addressing privacy concerns and safeguarding against information leakage to ensure robustness and trustworthiness in AI-powered systems.
@@ 5.3.2 ToolLLM and ChatDB as a multi active core-agent system
As mentioned in Section 4.4.2, incorporating multiple active core-agents within one system could pose challenges. To evaluate this scenario, two architectural variants were explored: LA2-A and LA2-B. In LA2-A, Figure 16, both ToolLLM and ChatDB retained their individual functionalities as distinct active core-agents. This case requires synchronization between the two active core-agents and opting for a consensus algorithm like Raft [70] would be a suitable choice. To further optimize the memory footprint of the system, there will be one unique instance of Llama 3.1 8B shared between the two active core-agents and each one of them must inject the adequate profile dynamically either as a pluggable trained module like LoRA or using a system prompt.
@@ 5.3.4 Hybrid multi active/passive core-agents system
Nonetheless, it is crucial to address the security vulnerability uncovered through the application of LLM-Agent-UMF within LA4’s design. Indeed, the external API calling mechanisms utilized by Toolformer and Confucius passive core-agents are neither monitored nor safeguarded against potential information leakage. Following the guidance provided by LLM-Agent-UMF, this issue can be resolved either by implementing a dedicated security module on each of these passive core-agents or as supplementary measures within the active core-agent’s security module, thus centralizing the management of information safeguarding processes. The optimal selection of technologies and implementation details will be explored in future work.
@@ 6 Conclusion and future work
In this paper, we introduce a structural component within LLM-based agents named the “core-agent”. This component is engineered to address the architectural ambiguities that software developers encounter and foster better understanding of the interacting entities within LLM-powered agents. We propose the LLM-Agent-UMF, a comprehensive framework, developed and evaluated using the ATRAF’s Architectural Framework Tradeoff and Risk Analysis Method (AFTRAM) [16, 17], for modeling the structure of the agent, explicating each of core-agent’s five modules: planning, memory, profile, action, and security. Subsequently, we classified core-agents into passive and active categories and highlighted their structural and functional differences. Based on this classification, we designed uniform and hybrid multi-core agent architectures. Most prominently, the one-active-many-passive architecture exploits the full potential of both active and passive core-agents, striking a balance between easiness of development and the power of hybrid architectures. By applying our framework to state-of-the-art agents, we identified within their structures core-agents and their constituting internal modules, which assisted us in the classification process. This allowed us to recognize the individual characteristics of each agent, discern their limitations, and discover potential prospects of merging different functionalities into a single multi-core agent.
Our work prepares the foundation for the development of LLM-based agents with a clear delineated structure that leverages the power of core-agents. The progressive adoption of LLM-Agent-UMF will further attest to its effectiveness. An interesting future direction to improve it involves finding solutions to simplify the implementation of multi-active core-agent architectures. In fact, they suffer from challenges related to synchronization that necessitate further investigation. In this context, we see two promising avenues: Integrating a consensus algorithm like Raft to elect a leader responsible for the coordination; Otherwise, integrating a central gateway that is solely responsible of selecting the most suitable active core-agent to handle the user request based on factors like load, availability, and domain. Each active core-agent shall register with the gateway, providing information about their capabilities and status. The selected core-agent processes the task and sends the response back to the user through the gateway.